OpenStudy (anonymous):

Solve 3x2 + 4x = 2 with the quadratic equation! I got -2-+ sqr of 40/3 but I don't see it in the answer choices. help?

OpenStudy (anonymous):

quantity of 2 plus or minus square root of 10 all over 6 quantity of negative 2 plus or minus 2 square root of 10 all over 3 quantity of negative 2 plus or minus square root of 10 all over 3 quantity of 4 plus or minus square root of 10 all over 3 these are the answer choices.

OpenStudy (anonymous):

for a quadratic eqn \[Ax^2 + Bx + C = 0\] the roots are \[x1 = \frac{ -B + \sqrt( B^2 -4AC) }{ 2A } and x2 = \frac{ -B - \sqrt(B^2-4AC) }{2A }\]

OpenStudy (anonymous):

try to solve quadratic equation with above method , u will surely get an answer

OpenStudy (anonymous):

can you draw it out cause its hard to understand it this way!!

OpenStudy (anonymous):

equation given in the ques is 3x2+4x=2 second step would be 3x2+4x−2=0 comparing above with general equation u will get A = 3 , B = 4 and C = -2 now just put the values of A,B and C in x1 and x2

OpenStudy (anonymous):

I got that and I put -4-+ sqr 4^2-4(3)(-2)/2(3)

OpenStudy (anonymous):

just simplify it more u will surely get your answer

OpenStudy (texaschic101):

-4 +- (sq rt -4^2 - 4(3)(-2))/2(3) -4 +- (sq rt 16 + 24)/6 -4 +- (sq rt 30)/6 I don't know...I got this

OpenStudy (anonymous):

idk how can you help?

OpenStudy (anonymous):

you made a mistake 16 +24 is 40 not 30

OpenStudy (texaschic101):

oops...your right

OpenStudy (anonymous):

@texaschic 101 u have done it correctly u just hve dne wrong addition addition should be 24 + 16 = 40

OpenStudy (anonymous):

yeah idk what to do afterwards

OpenStudy (anonymous):

I got the same thing but how is there a 10?

OpenStudy (anonymous):

can we write sqrt(40) as sqrt( 2 * 2 * 10) ?

OpenStudy (anonymous):

yeah I guess

OpenStudy (texaschic101):

oh...yes we can

OpenStudy (anonymous):

can we write sqrt(2*2*10) as 2 sqrt (10) ?

OpenStudy (texaschic101):

sq rt 4 * sq rt 10 = 2 sq rt 10

OpenStudy (anonymous):

oh my gosh it makes sense now thanj you so much!!!!!!!!!!!!!!!!

OpenStudy (texaschic101):

yes thanks...I am so not good at quadratic formula stuff

OpenStudy (texaschic101):

so the answer is b because we divided by 2...right ?

OpenStudy (anonymous):

I thought its c..

OpenStudy (anonymous):

no wait it is b..right?

OpenStudy (anonymous):

you are right @ninashabrang

OpenStudy (texaschic101):

your right....I didn't see the 2 in there

OpenStudy (anonymous):

oh gosh..okay ahha thanks again

OpenStudy (anonymous):

wait. so its c?

OpenStudy (anonymous):

yeah it should be c

OpenStudy (anonymous):

okay!